OpenStudy (anonymous):

Which polynomial is prime? 4x2 - 17x + 1 4x2 + 13x 8x2 + 7x - 16 x2 - 4

OpenStudy (anonymous):

I need the answer asap! Thanks to anyone who can help!

OpenStudy (anonymous):

\[\Large x^2-4=x^2-2^2=(x-2)(x+2)\] \[\Large 4x^2 + 13x=x(4x+13)\] These I can't factor them out... \[\Large 4x^2 - 17x + 1\] \[\Large 8x^2 + 7x - 16\]

OpenStudy (anonymous):

I don't understand well enough the question, if this is what you mean...